A light bulb is at the bottom of a swimming pool of 20 m. Calculate the diameter of the disc through which light emerges at the surface.
1
Expert's answer
2019-01-16T10:25:12-0500

h=20 m.

n=1.33

The angle of incidence:

r=90°

μ =sin⁡r/sini

1.33 =( sin 90°)/sini

i =sin^(-1)⁡〖(1/1.33) 〗 = 48.75°

D=2(20)tan 48.75°=45.6 m.